# High-OAM Deep Ultraviolet Twisted Light Generation for RF-Photoinjector Applicationsthanks: This document is the results of the research project funded by the Russian Science Foundation (Project No. 23-62-10026) [RSF]

### v3

@@ Abstract
We report on the generation and characterization of ultraviolet (the wavelength of 266 nm) twisted light with a high orbital angular momentum (OAM) using three types of fabricated diffractive optical elements (DOEs): a reflective fork grating, a high-charge spiral phase plate (SPP), and binary axicons. All elements were integrated into a drive-laser beamline of an electron RF-photoinjector, enabling direct evaluation under accelerator-relevant conditions.
The SPP produced a high-purity Laguerre–Gaussian mode with the OAM $\ell=64$ and a measured conversion efficiency of  80%. Binary axicons generated quasi–Bessel twisted light with the topological charges up to m = 10, exhibiting low divergence and stable multi-lobe ring structures. The fork grating reliably produced lower-order modes, $\ell=2$–8, with a good agreement between simulations and cylindrical-lens diagnostics.
These results constitute, to our knowledge, the first comprehensive experimental demonstration of deep-UV high-OAM beams generated with fabricated DOEs and validated through mode-conversion measurements. The demonstrated techniques are compatible with high-power UV laser systems used in RF-photoinjectors and offer a practical route toward structured photocathode illumination and the generation of relativistic vortex electron at a particle accelerator facility.
@@ 1 Introduction
Optical beams of the so-called twisted light [Allen92, AndrewsBabiker12, Franke-Arnold08, Bliokh17, Babiker19, Floettmann20, Karlovets21, Ivanov22, Fu24], carrying orbital angular momentum (OAM) play an increasingly important role in modern photonics, enabling advanced control of light–matter interactions [Schimmoller24], high-dimensional quantum communication protocols [Brandt20, Fickler20, Cao20], and structured-beam excitation in electron and plasma systems [Denoeud17]. Generating well-defined OAM modes in the deep ultraviolet (UV) range is particularly relevant for photocathode-based electron sources [Bliokh17, Floettmann20, Karlovets21, Ivanov22, Pavlov24] and nonlinear UV photonics, where high spatial coherence, low divergence, and precise transverse-mode control are required.
@@
However, producing stable high-charge OAM beams in the deep UV remains technically challenging. Spatial light modulators and digital micromirror devices, widely used at visible wavelengths, suffer from strong absorption, laser-induced damage, and low diffraction efficiency in the UV range. Binary spiral zone plates (SZPs) operate in this spectral range but are intrinsically limited to about 40% conversion efficiency [Dyatlov26]. Classical fork gratings [Solomonov24] can offer higher efficiency, yet they restrict the accessible OAM values to integer multiples of the diffraction order and often require tight focusing. Spiral phase plates (SPPs), while conceptually ideal for generating arbitrary OAM values [Khonina20], demand nanometer-scale surface accuracy over millimeter-sized apertures [Oemrawsingh04] and must withstand significant UV fluence. Binary axicons, widely used to generate Bessel-like twisted light (vortex) [Osintseva22, Knyazev23, Bazdyrev25], provide robustness and tunability but introduce characteristic ring fragmentation due to their discretized phase profiles. As a consequence, none of these approaches alone provides a universal, high-fidelity, UV-compatible solution for high-order OAM generation.
@@
In this work, we demonstrate ultraviolet twisted light generation using three fabricated diffractive optical elements (DOEs): reflective fork gratings, high-charge spiral phase plates, and binary axicons, all integrated into a drive-laser system of an RF-photoinjector [Gacheva14]. We generate Laguerre–Gaussian (LG) modes with the OAM values up to $\ell=64\hbar$, verify their modal structure using cylindrical-lens mode conversion, and compare the measured intensity distributions with numerical simulations. The generated beams exhibit high spatial fidelity, stable operation under UV irradiation, and low divergence suitable for photocathode illumination. The fabrication techniques used here rely on robust, scalable processes compatible with fused silica, SU-8 resists, and metal-coated substrates, enabling durable large-aperture UV optics.
@@
To the best of our knowledge, this is the first comprehensive experimental demonstration of the deep-UV high-OAM beams produced using fabricated DOEs and validated through the mode-conversion diagnostics. These results are directly relevant for the next-generation accelerator facilities. Relativistic vortex electrons are attracting increasing interest in nuclear and hadronic physics, where OAM offers a new degree of freedom for probing spin-structure phenomena and may complement the conventional spin-polarized beams [Floettmann20, Ivanov22, SchattschneiderPRB12, Wu19, An25]. Structured photocathode illumination is likewise becoming important for accelerator-based light sources, including polarized synchrotron-radiation facilities, free-electron lasers, and Compton $\gamma$-ray sources, all of which benefit from precise control of the transverse phase space of the injected electrons. The demonstrated deep-UV high-OAM beam generation thus can provide [RSF] a practical route toward the realization of MeV-scale vortex electrons in modern high-brightness photoinjectors.
@@ 2.1 A Reflective Diffraction Fork Grating with a Topological Charge m = 2
A fork grating introduces a phase dislocation of topological charge $m$, generating optical vortices in the diffracted orders. The phase profile employed in this work follows Ref. [Allen92]:
@@
In this context, the topological charge of the grating $m$ is the number of $2\pi$ phase windings embedded into the grating period. When a Gaussian beam is diffracted, each diffraction order acquires an optical vortex with topological charge of the beam [Kotlyar22]
@@
The resulting binary mask is shown in Fig. 2. The pattern was transferred onto an aluminum mirror via laser engraving, yielding a reflective grating with a $1\times 1~\mathrm{mm}$ active area and a spatial resolution of approximately $30~\mu\mathrm{m}$. A microscope image of the fabricated device is presented in Fig. 2.
Figure 1: A binarized fork-grating mask.
Figure 2: A microscope image of the fork grating.
@@ 2.2 A Spiral Phase Plate with a Topological Charge m = 64
A spiral phase plate (SPP) generates an azimuthally varying phase through a helical surface relief. The phase shift introduced by the plate is given by [Oemrawsingh04]
@@
For UV operation at $\lambda=266~\mathrm{nm}$ using fused silica ($n=1.49$), the required total height variation is approximately $h_{s}\approx 34~\mu\mathrm{m}$. Implementing a topological charge of $m=64$ is achieved by dividing the spiral into 64 discrete azimuthal sectors, each introducing a $2\pi$ phase step. This corresponds to a height increment of about $532~\mathrm{nm}$ per sector. The resulting two-dimensional phase topology is shown in Fig. 4.
The SPP was fabricated in fused silica using multi-level photolithography followed by high-precision plasma etching (Holo/Or, Israel). This manufacturing process is commonly employed for diffractive optical elements and provides the required micron-scale surface relief and sub-micron accuracy of the phase profile, ensuring stable performance under high-intensity UV illumination [Goebel96, Wang97, Kostyuk22, Shi24].
@@
2.3 Axicons with the Topological Charge m = 3 and 10
Bessel beams were modeled using the Wave-ThruMasks MATLAB package [Kameshkov19], which computes field propagation through transparent phase elements within the scalar diffraction theory using the Rayleigh–Sommerfeld integral [Goodman05]. The binary phase distribution used to generate twisted light Bessel beams is
@@ 2.3 Axicons with the Topological Charge m = 3 and 10
where $m$ is the topological charge, $\varphi$ the azimuthal angle, $k$ the radial wave number, and $r=\sqrt{x^{2}+y^{2}}$. An example mask for $m=3$ is shown in Fig. 5(a).
@@
The encoded topology was first transferred onto a glass photomask in an iron-oxide layer. The axicons were then fabricated via photolithography in SU-8 photoresist deposited on leucosapphire substrates, which provide high ultraviolet transmittance. The polymer crosslinked network of SU-8 also offers high resistance to ionizing radiation [Nazmov04]. Prior to resist deposition, the substrates were cleaned sequentially in acetone and in a 5% aqueous sulfuric-acid solution. To achieve the required film thickness, SU-8 2002 resist was diluted with SU-8 2000 Thinner.
@@
3 Numerical Modeling of the Generated Beams
Numerical simulation of the fork grating and the SPP was performed using a standard angular-spectrum propagation framework [Goodman05]. The incident field was taken to be a Gaussian beam,
$E_{\mathrm{G}}(x,y)=E_{0}\exp\left[-\frac{x^{2}+y^{2}}{w_{0}^{2}}\right],$ (10)
where $x$ and $y$ are the transverse coordinates, $w_{0}$ is the initial beam waist, and $E_{0}$ is the input field amplitude. This field was multiplied by the corresponding DOE phase: $\Phi(x,y)$ from Eq. 1 for the fork grating and $\varphi_{0}(\theta)$ from Eq. 5 for the SPP:
$E_{\mathrm{fork}}(x,y)=E_{G}(x,y)\exp\left[i\Phi(x,y)\right],$ (11)
$E_{\mathrm{SPP}}=E_{G}(x,y)\exp\left[i\varphi_{0}(\theta)\right]A(R).$ (12)
To suppress boundary artifacts and mitigate residual on-axis intensity that appear in high-charge modes, an apodization function was introduced:
$A(R)=\exp\left[\left(\frac{R}{R_{0}}\right)^{p}\right]\cdot\left(1-\exp\left[\left(\frac{R}{R_{C}}\right)^{q}\right]\right),$ (13)
where $R=\sqrt{x^{2}+y^{2}}$, $R_{0}$ defines the outer truncation radius, and $R_{C}$ specifies the central suppression zone.
Free-space propagation was computed using the angular-spectrum method,
$E(x,y,z)=\mathcal{F}^{-1}\left\{\mathcal{F}[E(x,y)]H_{prop}(f_{x},f_{y},z),\right\}$ (14)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the forward and inverse 2D Fourier transforms, $f_{x}$ and $f_{y}$ are the spatial frequencies, and the transfer function is
$H_{\mathrm{prop}}(f_{x},f_{y},z)=\exp\left[-i\pi\lambda z(f_{x}^{2}+f_{y}^{2})\right],$ (15)
When a lens was included, its quadratic phase was applied in the Fourier domain, after which the field was propagated to the observation plane using the same operator. The resulting optical intensity distributions were computed as
$I(x,y)=|E_{\mathrm{out}}(x,y)|^{2},\qquad I_{\mathrm{norm}}(x,y)=\frac{I(x,y)}{I_{\mathrm{max}}(x,y)}.$ (16)
@@ 3 Numerical Modeling of the Generated Beams
Figure 8: Simulated quasi–Bessel beams generated by binary axicons with topological charges $m=3$ and 1010 (left to right). Each pattern exhibits an $m$-lobe intensity ring, while the binary phase structure leads to a segmented rather than continuous ring profile.
@@
Similarly to the Laguerre–Gaussian simulations, quasi–Bessel beams produced in the focal region of the fabricated axicons were numerically modeled following the methodology of Ref. [Bazdyrev25]. The results are shown in Fig. 8. For all topological charge values, the characteristic ring-shaped intensity distribution is reproduced, consisting of $m$ localized bright lobes whose number corresponds to the topological charge of each axicon. Small parasitic background features appear as weak spiral-like artifacts.
@@
We also note that the binary phase pattern can lead to an effective interference of several overlapping wave components carrying slightly different OAMs. In this interpretation, the quasi–Bessel field can be viewed not as a single-mode vortex, but as a superposition of nearby OAM states [Vasilyeu09, Kovalev15, Saadati-Sharafeh20]. The average topological charge would still correspond to the design value $m$, while the OAM can exhibit a finite dispersion, defining an effective OAM bandwidth. This qualitative picture is consistent with the observed fragmented ring structure and is commonly discussed for binary axicons.
@@
In contrast, an ideal (non-binary) axicon produces a continuous, smooth ring [Osintseva22]. In the present case, the discretization inherent to the binary phase patterns results in a fragmented ring composed of distinct intensity maxima, consistent with the expected behavior of binary axicons.

### v4

# High-OAM Deep Ultraviolet Twisted Light Generation for RF-Photoinjector Applicationsthanks: This document is the results of the research project funded by the Russian Science Foundation (Project No. 23-62-10026) [37]

@@ Abstract
The generation of relativistic vortex electron beams via photoemission requires ultraviolet laser beams with well-controlled orbital angular momentum (OAM) and compatibility with radio-frequency (RF) photoinjector drive-laser systems. High-OAM vortex beams at a wavelength of 266 nm are generated using three fabricated diffractive optical elements integrated into an operational photoinjector beamline: a reflective fork grating, a high–topological-charge spiral phase plate, and binary axicons. The spiral phase plate produces a high-purity Laguerre–Gaussian mode with an OAM of $\ell=64\hbar$ and a conversion efficiency of 80%, whereas binary axicons generate low-divergence quasi–Bessel beams forming a superposition of multiple OAM states with a finite OAM bandwidth imposed by their binary phase structure. Fork gratings provide flexible access to lower OAM values and enable robust modal diagnostics. The generated beams are characterized using cylindrical-lens mode conversion and radial intensity analysis, demonstrating practical control of both the OAM content and spectral bandwidth of ultraviolet structured light for accelerator-based applications.
@@ 1 Introduction
Optical beams of the so-called twisted light [1, 3, 16, 7, 4, 15, 24, 22, 17], carrying orbital angular momentum (OAM) play an increasingly important role in modern photonics, enabling advanced control of light–matter interactions [42], high-dimensional quantum communication protocols [8, 14, 9], and structured-beam excitation in electron and plasma systems [12]. Generating well-defined OAM modes in the deep ultraviolet (UV) range is particularly relevant for photocathode-based electron sources [7, 15, 24, 22, 36] and nonlinear UV photonics, where high spatial coherence, low divergence, and precise transverse-mode control are required.
@@
However, producing stable high-charge OAM beams in the deep UV remains technically challenging. Spatial light modulators and digital micromirror devices, widely used at visible wavelengths, suffer from strong absorption, laser-induced damage, and low diffraction efficiency in the UV range. Binary spiral zone plates (SZPs) operate in this spectral range but are intrinsically limited to about 40% conversion efficiency [13]. Classical fork gratings [44] can offer higher efficiency, yet they restrict the accessible OAM values to integer multiples of the diffraction order and often require tight focusing. Spiral phase plates (SPPs), while conceptually ideal for generating arbitrary OAM values [25], demand nanometer-scale surface accuracy over millimeter-sized apertures [34] and must withstand significant UV fluence. Binary axicons, widely used to generate Bessel-like twisted light (vortex) [35, 26, 5], provide robustness and tunability but introduce characteristic ring fragmentation due to their discretized phase profiles. As a consequence, none of these approaches alone provides a universal, high-fidelity, UV-compatible solution for high-order OAM generation.
@@
In this work, we demonstrate the generation of deep-ultraviolet twisted light using three fabricated diffractive optical elements (DOEs): reflective fork gratings, high topological charge spiral phase plate, and binary axicons, all integrated into a drive-laser system of an RF-photoinjector [19]. We generate Laguerre–Gaussian (LG) modes with the OAM values up to $\ell=64\hbar$, verify their modal structure using cylindrical-lens mode conversion, and compare the measured intensity distributions with numerical simulations In contrast, binary axicons produce low-divergence quasi–Bessel beams whose OAM content forms a superposition of multiple topological charges, reflecting the binary radial phase modulation. The generated beams exhibit high spatial fidelity, stable operation under UV irradiation, and low divergence suitable for photocathode illumination. The fabrication techniques used here rely on robust, scalable processes compatible with fused silica, SU-8 resists, and metal-coated substrates, enabling durable large-aperture UV optics.
@@
To the best of our knowledge, this is the first comprehensive experimental demonstration of the deep-UV high-OAM beams produced using several types of fabricated DOEs and validated through the mode-conversion diagnostics. These results are directly relevant for the next-generation accelerator facilities with the so-called vortex electrons - massive counterparts of twisted photons. Relativistic vortex electrons are attracting increasing interest in nuclear and hadronic physics, where OAM offers a new degree of freedom for probing spin-structure phenomena and may complement the conventional spin-polarized beams [15, 22, 40, 50, 2]. Structured photocathode illumination is likewise becoming important for accelerator-based light sources, including polarized synchrotron-radiation facilities, free-electron lasers, and Compton $\gamma$-ray sources, all of which benefit from precise control of the transverse phase space of the injected electrons. The demonstrated deep-UV high-OAM beam generation thus can provide [37] a practical route toward the realization of MeV-scale vortex electrons in modern high-brightness photoinjectors.
@@ 2.1 A reflective diffraction fork grating with a topological charge m = 2
A fork grating introduces a phase dislocation of topological charge $m$, generating optical vortices in the diffracted orders. The phase profile employed in this work follows Ref. [1]:
@@
In this context, the topological charge of the grating $m$ is the number of $2\pi$ phase windings embedded into the grating period. When a Gaussian beam is diffracted, each diffraction order acquires an optical vortex with topological charge of the beam [28]
@@
The resulting binary mask is shown in Figure 2. The pattern was transferred onto an aluminum mirror via laser engraving, yielding a reflective grating with a $1\times 1~\mathrm{mm}$ active area and a spatial resolution of approximately $30~\mu\mathrm{m}$. A microscope image of the fabricated device is presented in Figure 2.
Figure 1: A Binarized fork-grating mask generated according to Eq. 1. The grayscale corresponds to binary transmittance values (0 or 1). Coordinates are shown in normalized units.
Figure 2: A microscope image of the reflective fork grating laser-engraved on an aluminum-mirror substrate. The fabricated structure reproduces the binary mask of Figure 2. The active area is $1\times 1~\mathrm{mm}$ with a spatial resolution of $30~\mu\mathrm{m}$.
@@ 2.2 A spiral phase plate with a topological charge m = 64
A spiral phase plate (SPP) generates an azimuthally varying phase through a helical surface relief. The phase shift introduced by the plate is given by [34]
@@
For UV operation at $\lambda=266~\mathrm{nm}$ using fused silica ($n=1.49$), the required total height variation is approximately $h_{s}\approx 34~\mu\mathrm{m}$. Implementing a topological charge of $m=64$ is achieved by dividing the spiral into 64 discrete azimuthal sectors, each introducing a $2\pi$ phase step. This corresponds to a height increment of about $532~\mathrm{nm}$ per sector. The resulting two-dimensional phase topology is shown in Figure 4.
The SPP was fabricated in fused silica using multi-level photolithography followed by high-precision plasma etching (Holo/Or, Israel). This manufacturing process is commonly employed for diffractive optical elements and provides the required micron-scale surface relief and sub-micron accuracy of the phase profile, ensuring stable performance under high-intensity UV illumination [20, 48, 27, 43].
@@
2.3 Binary axicons with the topological charge m = 3 and 10
Bessel beams were modeled using the Wave-ThruMasks MATLAB package [23], which computes field propagation through transparent phase elements within the scalar diffraction theory using the Rayleigh–Sommerfeld integral [21]. The binary phase distribution used to generate twisted light Bessel beams is
@@ 2.3 Binary axicons with the topological charge m = 3 and 10
where $m$ is the topological charge, $\varphi$ the azimuthal angle, $k$ the radial wave number, and $r=\sqrt{x^{2}+y^{2}}$. An example mask for $m=3$ is shown in Figure 5(a).
@@
The encoded topology was first transferred onto a glass photomask in an iron-oxide layer. The axicons were then fabricated via photolithography in SU-8 photoresist deposited on leucosapphire substrates, which provide high ultraviolet transmittance. The polymer crosslinked network of SU-8 also offers high resistance to ionizing radiation [33]. Prior to resist deposition, the substrates were cleaned sequentially in acetone and in a 5% aqueous sulfuric-acid solution. To achieve the required film thickness, SU-8 2002 resist was diluted with an SU-8 2000 thinner.
@@
3 Numerical simulation of the generated beams
Numerical simulation of the fork grating and the SPP was performed using a standard angular-spectrum propagation framework [21]. A Gaussian input beam was assumed and multiplied by the corresponding phase profile of the diffractive optical element: $\Phi(x,y)$ for the fork grating (Eq. 1) and $\varphi_{0}(\theta)$ for the SPP (Eq. 5). Free-space propagation and lens transformations were implemented using Fourier-domain transfer operators.
To suppress boundary effects and residual on-axis intensity that become pronounced for high-OAM modes, a radial apodization was applied, introducing an outer truncation and a central suppression region. The specific functional form was chosen to minimize numerical artifacts without affecting the intrinsic transverse mode structure.
@@ 3 Numerical simulation of the generated beams
Figure 8: Simulated quasi–Bessel beams generated by binary axicons with topological charges $m=3$ and $10$ (left to right). Each pattern exhibits an $m$-lobe intensity ring, while the binary phase structure leads to a segmented rather than continuous ring profile.
@@
Similarly to the Laguerre–Gaussian simulations, quasi–Bessel beams produced in the focal region of the fabricated axicons were numerically modeled following the methodology of Ref. [5]. The results are shown in Figure 8. For all topological charge values, the characteristic ring-shaped intensity distribution is reproduced, consisting of mm localized bright lobes whose number corresponds to the topological charge of each axicon. Small parasitic background features appear as weak spiral-like artifacts.
@@
We also note that the binary phase pattern can lead to an effective interference of several overlapping wave components carrying slightly different OAMs. In this interpretation, the quasi–Bessel field can be viewed not as a single-mode vortex, but as a superposition of nearby OAM states [46, 30, 38]. The average topological charge would still correspond to the design value $m$, while the OAM can exhibit a finite dispersion [10], defining an effective OAM bandwidth. This qualitative picture is consistent with the observed fragmented ring structure and is commonly discussed for binary axicons.
@@
In contrast, an ideal (non-binary) axicon produces a continuous, smooth ring [35]. In the present case, the discretization inherent to the binary phase patterns results in a fragmented ring composed of distinct intensity maxima, consistent with the expected behavior of binary axicons.
3.1 OAM spectrum decomposition
To quantify the OAM content of the generated fields, we project the numerically obtained complex transverse field onto azimuthal harmonics $\exp(i\ell\varphi)$ [45, 47]. The projection coefficient is evaluated as
$c_{\ell}\propto\int E(r,\varphi)\,\exp(-i\ell\varphi)\,W(r)\,r\,\mathrm{d}r\,\mathrm{d}\varphi,$ (10)
where $W(r)$ accounts for the finite aperture. The normalized spectrum is defined as
$P_{\ell}=|c_{\ell}|^{2}\qquad\sum_{\ell}P_{\ell}=1.$ (11)
For Laguerre–Gaussian-like beams generated by fork gratings and spiral phase plates, the field contains a single azimuthal phase factor $\exp(i\ell_{0}\varphi)$. Owing to the orthogonality of azimuthal harmonics, only $\ell=\ell_{0}$ survives upon integration, yielding a narrow single-peaked OAM spectrum (Figure 9(a–b)). In fork gratings, different diffraction orders carry distinct, well-defined topological charges, producing separated peaks at $\ell=\pm 2$, $\pm 4$, and $\pm 8$.
In contrast, binary axicons generate quasi–Bessel beams with coupled radial and azimuthal structures. The field can be represented as a superposition of opposite OAM components $\pm\ell$ whose weights vary with radius [39, 18]. As a result, the azimuthal projection yields multiple non-vanishing components and a broadened OAM spectrum with a finite intrinsic bandwidth (Figure 9(c–d)).
To reduce numerical cancellation caused by rapid radial phase oscillations, the aperture is divided into $N$ concentric annuli, within which local projections $c_{\ell}^{(j)}$ are evaluated and summed as
$P_{\ell}=\sum_{j}|c_{\ell}^{(j)}|^{2}.$ (12)
This procedure preserves the physical OAM content while improving numerical stability.
Figure 9: Orbital angular momentum (OAM) spectrum of the generated beams obtained by numerical projection onto azimuthal harmonics. (a) Fork grating: discrete, well-separated peaks at $\ell=\pm 2$, $\pm 4$, and $\pm 8$, corresponding to the individual diffraction orders of the grating and indicating high OAM spectral purity. (b) Spiral phase plate (SPP): a single narrow peak centered at the design topological charge, demonstrating the generation of a nearly pure OAM eigenstate. (c) Binary axicon with topological order $m=3$ and (d) binary axicon with $m=10$: broadened and structured OAM spectra resulting from a superposition of multiple azimuthal harmonics with opposite topological charges, characteristic of quasi–Bessel beams.
3.2 Control of mode purity in beams generated by a binary axicons
As shown in Figure 9(a–b), fork gratings and spiral phase plates generate nearly pure OAM eigenstates. In contrast, binary spiral axicons produce coherent superpositions of azimuthal harmonics, leading to intrinsically mixed OAM spectrum.
[…]
Variation of the thickness $h$ modifies the effective complex transmission function of the axicon and leads to the appearance of a non-negligible $\ell=0$ contribution in the OAM spectrum (Figure 11). Even changes of several tens of nanometers noticeably reduce the modal purity. This behavior originates from deviations of the phase modulation from its optimal value combined with amplitude modulation introduced by the finite conductivity of the metal. In the experiment, such a contribution would manifest itself as a Gaussian-like on-axis intensity core superimposed on the annular Bessel structure.
Figure 11: Projected OAM spectra $|c_{\ell}|^{2}$ of the field generated by the binary spiral axicon for different SU-8 thicknesses: (a) $h=180$ nm, (b) $h=200$ nm, and (c) $h=220$ nm. Variation of the phase modulation depth leads to the appearance of a non-negligible $\ell=0$ component and a redistribution of the relative weights of the $\pm\ell$ harmonics. All other parameters are kept fixed.